# Characterizing the MrDeepFakes Sexual Deepfake Marketplace

### 14 Oct 2024 v1

@@ 5.2 Creators and Sellers
Over 1,800 users have uploaded deepfake video content on MrDeepFakes, though we note that $657$ users (0.1%) uploaded over 95% of the site’s video content, inline with the distribution of user participation on other social media platforms [39]. Around 228 users have responded to requests for free or paid deepfake videos, of which 55 are “verified” sellers (Section 5.2.2); however, this is likely an underestimate of the total number of users creating videos since a significant fraction of communication occurs over private channels. 94.7% of paid and 72.9% of free requests are replied to and the median time to receive a response are 16 hours and 5 days, respectively.
Sellers responded to an average eight posts, but this is heavily skewed with a handful 10–15 sellers responding to an outsized number of requests. One seller responded to 531 requests alone. In these cases, sellers appear to aggressively respond to most posts with generic replies like “PM sent,” “Please check your PM, thank you,” and “Here is a sample of my work: […] dm me on discord : […]”
@@ 5.2.1 Marketing and Differentiation
Sellers commonly showcase their abilities to potential buyers by posting their own deepfake videos on MrDeepFakes’s tube site. In paid request threads, 365 (18.9%) of deepfake creators’ replies included links to deepfake videos on the tube site as well as embedded screenshots or thumbnails from past work. Because deepfake approaches are rapidly evolving and creators want to showcase only their best work, this also appears to lead to churn in hosted videos. We observed that $1948$ of $42986$ video URLs were inaccessible; 75.1% of deleted videos indicated that they had been deleted upon their creator’s request. Anecdotally, we find indications of increasing specialization among deepfake creators, like we observed with creators dedicated to deepfaking Guatemalan newscasters (Section 4.2). In response to a request seeking “somebody who would specialize [in] classic (1970s and 1980s)” deepfakes, a verified creator, moderator, and staff member responded:
@@ 5.2.2 Verified Creators
Deepfake creators can apply to become Verified Video Creators, who receive a “trusted” badge indicating that the user creates their own deepfake content and are able to monetize on the platform itself (e.g., paid video downloads, cryptocurrency payouts, advertising privileges in forum threads). Verified creators also bypass manual moderation of their content, which site administrators claim is a process that all uploads from unverified users undergo and is expected to take several days. Further, platform affordances incentivize verification: the tube site showcases verified creators’ videos, and a verified creator’s verification status appears as a colorful label alongside their forum posts. The application process aims to certify that they generate their own content. To demonstrate their required technical knowledge for deepfake creation, applicants must provide evidence that they possess datasets or images of applications they use; they must also describe their creation process. Beyond proving their capabilities, the applicant must also have reasonable standing within the MrDeepFakes community, as they are required to link to their tube site profile with at least three “active” videos (i.e., they have not been removed by moderators). Their account must also be e-mail verified, at least 30 days old, and cannot have received warnings from moderators.
@@ 5.3 Cost of a Custom Deepfake
The average price advertised by sellers is \$87.50; 35.1% of quoted prices are at least \$100, and 12.3% are at least \$200. The range in prices is significant: the most inexpensive advertised service was \$1 for a one-minute video with a 30-minute turnaround time with one condition: “No [blowjobs].” Conversely, the highest price discussed was \$$1500$. For 31 of the 58 request posts with explicit prices, the length of the video was included as a parameter; the length of a deepfake video is loosely correlated with the commission price ($R=0.29$). From the limited visibility of a public vantage point, we cannot meaningfully determine if prices have changed over time. However, we do find anecdotes from community leaders that describe how deepfake technology has shifted the economy of the broader and longer-standing “faking” market:
@@ 5.3.1 Factors Affecting Price
Several buyers noted the factors that affected the price that they were willing to pay; they were willing to pay more for videos with more than one target depicted, if the angles of the video are complicated, or if their target was male. Because women are predominantly targeted by sexual deepfakes, facesets and information regarding sex worker “body matches” for men are comparatively rarer. Several buyers described the factors they considered when choosing a seller to fulfill their request. Most commonly (12% of sampled requests), buyers explicitly stated that they sought high-quality deepfakes where a deepfake’s quality is typically defined by its resolution or the degree of achieved realism. Other factors included the estimated production turnaround time (1.9%) and the “professionalism” of the seller (0.8%). To proactively vet sellers, 12 requests asked interested creators to link to a portfolio of their past work; two other requests specified they were looking for a verified video creator (Section 5.2.2) or one with good reviews on the dedicated forum thread.
@@
The overall market for deepfakes has thus split into two tiers: (1) low-quality and fast “cheapfakes” and (2) high-end, custom “art.” The emergence of commodity deepfake applications (e.g., those marketed for face swapping, undressing, nudifying, etc. capabilities) [61, 44, 43] caters to the former, while seller scrutiny of cheapfake offers situates the paid requests marketplace on MrDeepFakes as offering the latter. These higher-end figures highlight the technical challenges posed in deepfake creation, which appear to require the expertise of a highly-skilled professional or the money to pay one.
@@
We find that this can occur even earlier between potential buyers and sellers than at the price negotiation step. Of the $32284$ videos that have non-null descriptions, using regex and keyword search, we identify that 24.9% of the video descriptions provide viewers with contact information or otherwise solicit private communication. This positions videos uploaded to the tube site as advertisements themselves. Offers to communicate via first-party direct messages on the MrDeepFakes site itself are found in 8.9% of videos. Third-party platforms are also popular; the most common third-party platforms for direct messages include Telegram, Discord, and e-mail.
@@ 5.3.2 Additional Revenue Streams
Verified creators on MrDeepFakes can accept donations, sell video downloads (versus online playback), and offer technical support, and provide other resources through MrDeepFakes tokens, a first-party digital currency where the lowest denomination of 100 tokens can be bought using \$1.99 worth of BTC, or directly through their Bitcoin address. Though we cannot measure how many users pay to download video content, we observe that the advertised download price varied greatly, ranging from \$2 to \$32 in tokens. In addition, video descriptions on the tube site commonly ($15039$ of $42986$ videos, 35%) included third-party links. About half (49%) are to payment sites or subscription signups on both mainstream (e.g., PayPal, Ko-Fi, Patreon) and adult content (e.g., OnlyFans, FansMine, Fan-Topia) platforms.
@@ 6 Deepfake Video Creation
In addition to serving deepfake video content and providing a marketplace for buyers and sellers, MrDeepFakes also hosts discussions about the creation of deepfake videos. Prior work has established that while nonexperts can successfully create deepfakes, the ability to create realistic deepfakes (e.g., not suspicious to a human or software detector) requires deep expertise using complex, difficult-to-use software [36]. An active community has emerged on MrDeepFakes to facilitate this process, using the forum to discuss best practices, offer tips and suggestions on improving deepfake quality, and provide detailed guides. In this section, we analyze forum discussion and the most popular guides for deepfake video production to document the requirements, challenges, and community solutions for the deepfake creation pipeline.
We specifically analyzed the “DeepFake Video Creation Tools” section of the forum, which is designated for members to discuss how to create deepfakes and also hosts guides and tutorials for deepfake creation, authored by expert community members, forum staff, administrators, moderators, and verified creators. The section contained 56.9% of all threads and 58.8% of all posts on the platform. We began by manually examining the most active guide threads (by number of replies), as well as guides linked in a pinned “quick start” post for newcomers. From there, we analyzed other linked threads, excluding topics unrelated to sexual deepfake videos, which uncovered 25 guide and tutorial threads containing $1229$ posts. We annotated discussions for troubleshooting topics or resource requirements for deepfake creation, surfacing three major themes: (1) access to facial image data, (2) software proficiency, and (3) powerful hardware. We discuss the challenges posed by each and how the MrDeepFakes community has developed solutions for each of them.
@@ 6.1 Facial Image Data
When facial image data is limited, users turn to the forum to request and share “facesets,” which are curated dumps of facial image data for a particular target. We find that facesets are often freely shared, as only two of the 711 threads that pertain to sharing facesets were labeled as paid solicitations. Of these threads, 81.6% contained links to third-party file sharing sites (Appendix B). In comparison, 326 threads request the creation of new facesets to target an individual; however, only 39.6% of these threads are replied to and only 8% are returned a link to a file sharing domain. Only three new faceset requests were paid offers. Experts’ discussion and purchase offers of facesets suggest that the curation of high quality facesets is labor-intensive and valuable:
@@ 6.2 Software
Most deepfakes on MrDeepFakes are created using DeepFaceLab (DFL) or FaceSwap (FS). DFL is preferred and lauded for its acceptance of “not safe for work” (NSFW) usage, technical capabilities, troubleshooting support, and lower system requirements:
@@
Both tools are free and open source but are complex to use. As a result, significant community effort is dedicated to streamlining and troubleshooting their usage, which we anecdotally find affects the adoption and popularity of deepfake software tools; troubleshooting comprised $3137$ (50.1%) of threads in the sexual deepfake video-related sections, representing 3,945 posts (41.6%) on the forum in aggregate. DFL is the subject of three pinned guides; the main DFL guide has the most views and replies (2M and 855, respectively), and the DFL repository creator and maintainer also appears to be an active member of the MrDeepFakes community, with an account adorned with a “Developer” badge. Several other authors of the DFL preprint [45] appear to be active members, with roles including developer, staff, administrator, and verified creator.
@@
Two “limited-data” tools stand out on MrDeepFakes: Roop/Rope, which have a moderator-written guide linked in the quick start, and SwapFace, which is the subject of the eighth-most active thread on the forum. In terms of output fidelity, some members praise the tools’ potential to produce high quality deepfakes for simple videos, while others lament their inability to do so for complicated videos. This parallels the striation of the deepfake market into demands for two categories: a lower-tier for “cheapfakes” and a higher-tier for complex scenes that require deeper expertise and extensive image data. The former is handled by these limited-data tools, whose output quality suffers for more complex tasks; buyers of the latter must then turn to the marketplace to pay an expert for their services.
@@
In our annotation of 141 linked Github repositories, we identified that 60 repositories (including forked repositories) were directly linked to 43 academic papers; in aggregate, these open source research code repositories were referenced in the forum over 400 times. The associated papers range in publication date: the oldest is a 2001 paper that achieves color correction between source and destination images using statistical analysis [51], and the six most recent papers were published in 2023, the year we collected our forum data [31, 49, 4, 2, 42, 64]. Of these papers, CVPR is by far the most prevalent venue, accounting for 13 of the 43 papers (30.2%); the remaining 30 papers are spread across 21 other venues, including well-known conferences ICCV, ECCV, NeurIPS, ICML, and SIGGRAPH. The frequent and up-to-date references to reputable academic publications demonstrates that the MrDeepFakes community follows cutting-edge research.
The community favorite, DFL (the Github repository was linked 203 times), the corresponding paper was initially published as an ArXiv preprint in 2020 and later published in the peer-reviewed journal Pattern Recognition in 2023. The delta between the two versions largely focuses on potential abuses of DFL, which both versions claim is most commonly used by VFX artists. In addressing the ethical issues of advancing deepfake technology, the authors “admit that deepfake techniques may affect public discourse quality and infringe upon the citizens’ portrait rights” when weaponized for “misinformation, manipulation, harassment, [or] persuasion.” Despite this, the authors argue that suppressing the publication would not stop its development and believe that it was their “responsibility to publish DeepFaceLab to the academia community formally.” Neither version directly addresses the use or abuse of deepfake technology to generate sexual imagery. In August 2024, the repository for DFL was disabled by GitHub for terms of service (ToS) violations.

### 02 Jan 2025 v2

@@ 5.3 Creators and Sellers
Over 1,800 users have uploaded deepfake video content on MrDeepFakes, though we note that $657$ users (0.1%) uploaded over 95% of the site’s video content, inline in line with the distribution of user participation on other social media platforms [nielsen2006participation]. Around 228 users have responded to requests for free or paid deepfake videos, of which 55 are “verified” sellers (Section 5.3.2); however, this is likely an underestimate of the total number of users creating videos since a significant fraction of communication occurs over private channels. 94.7% of paid and 72.9% of free requests are replied to, and the median time times to receive a response are 16 hours and 5 days, respectively.
Sellers responded to an average eight posts, but this is heavily skewed with a handful (10–15) of sellers responding to an outsized number of requests. One seller responded to 531 requests alone. In these cases, sellers appear to aggressively respond to most posts with generic replies like “PM sent,” “Please check your PM, thank you,” and Here is a sample of my work: […] dm me on discord: […]”
@@ 5.3.1 Marketing and Differentiation
Sellers commonly showcase their abilities to potential buyers by posting their own deepfake videos on MrDeepFakes’ s tube site. In paid request threads, 365 (18.9%) of deepfake creatorssellers’ replies included links to deepfake videos on the tube site as well as embedded screenshots or thumbnails from past work. Because deepfake approaches are rapidly evolving and creators want to showcase only their best work, this also appears to lead to churn in hosted videos. We observed that $1948$ of $42986$ video URLs were inaccessible; 75.1% of deleted videos indicated that they had been deleted upon their creator ’s creator request. Anecdotally, we find indications of increasing specialization among deepfake creators, like we observed with creators dedicated to deepfaking Guatemalan newscasters (Section 4.3). In response to a request seeking “somebody who would specialize [in] classic (1970s and 1980s)” deepfakes, a verified creator, moderator, and staff member responded:
@@
I definitely would, but my specialty is more placing today’s celebrities into 70s porn…If you suggest a celeb that inspires me, I might do it.
@@ 5.3.2 Verified Creators
Deepfake creators can apply to become Verified Video Creators, who receive a “trusted” badge indicating that the user creates their own deepfake content and are able to monetize on the platform itself (e.g., paid video downloads, cryptocurrency payouts, advertising privileges in forum threadsforum advertising privileges). Verified creators also bypass manual moderation of their content content moderation, which site administrators claim is a process that all uploads from unverified users undergo and is expected to take takes several days. Further, platform affordances incentivize verification: the tube site showcases verified creators’ videos, and a verified creator’s verification status appears as a colorful label alongside their forum posts. The application process aims to certify that they generate their own content. To demonstrate their required technical knowledge for deepfake creation, applicants must provide evidence that they possess datasets or images of applications they use; they must also describe their creation process. Beyond proving their capabilities, the applicant must also have reasonable standing within the MrDeepFakes community, as they are required to link to their tube site profile with at least three “active” videos (i.e., they have not been removed by moderators). Their account must also be e-mail verified, at least 30 days old, and cannot have received warnings from moderators.
@@ 5.4 Cost of a Custom Deepfake
The average price advertised by sellers is \$87.50; 35.1% of quoted prices are at least \$100, and 12.3% are at least \$200. The range in prices is significant: the most inexpensive least expensive advertised service was \$1 for a one-minute video with a 30-minute turnaround time with one condition: “No [blowjobs].” Conversely, the highest price discussed was \$$1500$. For 31 of the 58 request posts with explicit prices, the length of the video was included as a parameter; the length of a deepfake video is loosely correlated with the commission price ($R=0.29$). From the limited visibility of a public vantage point, we cannot meaningfully determine if prices have changed over time. However, we do find anecdotes from community leaders that describe how deepfake technology has shifted the economy of the broader and longer-standing “faking” market:
@@
Manual faking might have been killed by AI, but not faking in general. AI has made faking more accessible and now, it’s even harder to make money from it than before (it was already hard because of laws and policies on platforms like Patreon). Still, AI has sped up the faking process and lowered the barrier to entry for folks with less time or technical skills to make money off fakes (you aren’t limited to celeb fakes anymore, especially with tools like DeepFaceLab and Roop).
@@ 5.4.1 Factors Affecting Price
Several buyers noted the factors that affected the price that they were willing to pay; they were willing to pay more for videos with more than one target depicted, if the angles of the video are were complicated, or if their target was male. Because women are predominantly targeted by sexual deepfakespredominantly women are targeted by deepfake NCII, facesets and information regarding sex worker “body matches” for men are comparatively rarer. Several buyers described the factors they considered when choosing a seller to fulfill their request. Most commonly (12% of sampled requests), buyers explicitly stated that they sought high-quality deepfakes where a deepfake’s quality is typically defined by its resolution or the degree of achieved realism. Other factors included the estimated production turnaround time (1.9%) and the “professionalism” of the seller (0.8%). To proactively vet sellers, 12 requests asked interested creators to link to a portfolio of their past work; two other requests specified they were looking for a verified video creator (Section 5.3.2) or one with good reviews on the dedicated forum thread.
@@
Beware of cheap offers for \$10–50 fakes, they can’t possibly deliver quality. One guy who frequently advertises their services here did such a shit job for one of my clients that I would have banned him.
The overall market for deepfakes has thus split into two tiers: (1) low-quality and fast “cheapfakes” and (2) high-end, custom “art.” The emergence of commodity deepfake applications (e.g., those marketed for face swapping, undressing, nudifying, etc. capabilities) [westerlund2019deepfake, cbs2019cheapfake, paris2019cheapfake] caters to the former, while seller scrutiny of cheapfake offers situates the paid requests marketplace on MrDeepFakes as offering the latter. These higher-end figures highlight the technical challenges posed in deepfake creation, which appear to require the expertise of a highly-skilled professional or the money to pay one.
@@
The only answer is to find a legit SFX guy working in TV and film, pay him \$500 to \$1,500 at least. It’s probably even more to match the lighting and body parts nicely, it’ll take at least 2 weeks of work even for a professional.
@@
We find that this can occur even earlier between potential buyers and sellers than at sellers potentially establish private communication channels with potential buyers even before the price negotiation step. Of the $32284$ videos that have non-null descriptions, using regex and keyword search, we identify that 24.9% of the video descriptions provide viewers with contact information or otherwise solicit private communication. This positions videos uploaded to the tube site as advertisements themselves. Offers to communicate via first-party direct messages on the MrDeepFakes site itself are found in 8.9% of videos. Third-party platforms are also popular; the most common third-party platforms for direct messages include Telegram, Discord, and e-mail. We note limited use of third-party platforms that offer end-to-end encryption (E2EE) channels, identifying three references to Proton Mail addresses and 21 references to WhatsApp contact information. While Telegram offers E2EE in its “Secret Chat” mode, the default is not.
@@ 5.4.2 Additional Revenue Streams
Verified creators on MrDeepFakes can accept donations, sell video downloads (versus online playback), and offer technical support, and provide paid technical support or other resources through MrDeepFakes tokens, a first-party digital currency where the lowest denomination of 100 tokens can be bought using \$1.99 worth of BTC, or directly through their Bitcoin address. After inspecting the MrDeepFakes token purchase process, we found that token purchases occur only through a third-party cryptocurrency payment processor. While Bitcoin transactions are public, we would only know what address(es) to inspect if we purchased tokens ourselves, directly contributing to site revenue. Further, without making repeated token purchases over time, we would not be able to determine if transactions are split across different addresses. Therefore, we chose to not pursue further analysis of MrDeepFakes token purchase volumes. Though we cannot measure how many users pay to download video content, we observe that the advertised download price varied greatly, ranging from \$2 to \$32 in tokens.
Similarly, we collected and attempted analysis for 292 unique Bitcoin addresses provided in designated video metadata fields for direct donations. However, because of cryptocurrency transaction anonymity, we could not determine which transactions were related to deepfake services. Bounding transactions by associated MrDeepFakes accounts’ creation timestamps, we did find that 20% of addresses received at least \$1,000. Ultimately, we note that these estimates are too loosely bounded to make verified claims about deepfake-related revenue for these addresses. In addition, video descriptions on the tube site commonly ($15039$ of $42986$ videos, 35%) included third-party links. About half (49%) are to payment sites or subscription signups on both mainstream (e.g., PayPal, Ko-Fi, Patreon) and adult content (e.g., OnlyFans, FansMine, Fan-Topia) platforms.
@@ 6 Deepfake Video Creation
In addition to serving deepfake video content and providing a marketplace for buyers and sellers, MrDeepFakes also hosts discussions about the creation of deepfake videos. Prior work has established that while nonexperts can successfully create deepfakes, the ability to create realistic deepfakes (e.g., not suspicious to a human or software detector) requires deep expertise using complex, difficult-to-use software [mehta2023can]. An active community has emerged on MrDeepFakes to facilitate this process, using the forum to discuss best practices, offer tips and suggestions on improving deepfake quality, and provide detailed guides. In this section, we analyze forum discussion and the most popular guides for deepfake video production to document the requirements, challenges, and community solutions for the deepfake creation pipeline.
We specifically analyzed
6.1 Methodology
We specifically perform content analysis the “DeepFake Video Creation Tools” section of the forum, which is designated for members to discuss how to create deepfakes and also hosts guides and tutorials for deepfake creation, authored by expert community members, forum staff, administrators, moderators, and verified creators. The section contained 56.9% of all threads and 58.8% of all posts on the platform. We began by manually examining the most active guide threads (by number of replies), as well as guides linked in a pinned “quick start” post for newcomers. From there, we analyzed other linked threads, excluding topics unrelated to sexual deepfake videos, which uncovered 25 guide and tutorial threads containing $1229$ posts. We annotated discussions for troubleshooting topics or resource requirements for deepfake creation, surfacing three major themes: (1) access to facial image data, (2) software proficiency, and (3) powerful hardware. We discuss the challenges posed by each and how the MrDeepFakes community has developed solutions for each of them.
While we attempted large-scale, automated analysis of forum posts to understand the deepfake video creation process through BERTopic topic modeling, we encountered several obstacles. First, the topics that emerged did not provide insight beyond what the structure of the forum already offered (e.g., clusters about facesets, tools for deepfake video creation). Second, this approach fails to scale the process of revealing interesting thematic nuance within clusters, requiring manual annotation of topic clusters to effectively apply this method. Therefore, we ultimately pursued our manual approach.
@@ 6.2 Facial Image Data
Crowdsourced Facesets.When facial image data is limited, users turn to the forum to request and share “facesets,” which are curated dumps of facial image data for a particular target. We find that facesets are often freely shared, as only two of the 711 threads that pertain to sharing facesets were labeled as paid solicitations. Of these threads, 81.6% contained links to third-party file sharing sites (Appendix B). In comparison, 326 threads request the creation of new facesets to target an individual; however, only 39.6% of these threads are replied to and only 8% are returned a link to a file sharing domain. Only three new faceset requests were paid offers. Experts’ discussion and purchase offers of facesets suggest that the curation of high quality facesets is labor-intensive and valuable:
@@
I need someone to make me quality facesets…maybe 3 or 4 a month. I’ll pay you in crypto (enough to pay some if not all your rent, depending on where you live)
@@ 6.3 Software
Most deepfakes on MrDeepFakes are created using DeepFaceLab (DFL) or FaceSwap (FS). DFL is preferred and lauded by the community for its acceptance of “not safe for work” (NSFW) usage, technical capabilities, troubleshooting support, and lower system requirements:
@@
The FS dev team will not acknowledge any NSFW. The DFL dev has no issues with NSFW.
@@
DFL is far ahead in dealing with face obstructions. FS is working on it, but nothing has been released yet.
Both tools are free and open source but are complex to use. As a result, significant community effort is dedicated to streamlining and troubleshooting their usage, which we anecdotally find affects the adoption and popularity of deepfake software tools; troubleshooting comprised $3137$ (50.1%) of threads in the sexual deepfake video-related sections, representing 3,945 posts (41.6%) on the forum in aggregate. DFL is the subject of three pinned guides; the main DFL guide has the most views and replies (2M and 855, respectively), and the DFL repository creator and maintainer also appears to be an active member of the MrDeepFakes community, with an account adorned with a “Developer” badge. Several other authors of the DFL preprint [perov2021dfl] appear to be active members, with roles including developer, staff, administrator, and verified creator.
@@
Roop, Rope, and SwapFace.Two “limited-data” tools stand out on MrDeepFakes: Roop/Rope, which have a moderator-written guide linked in the quick start, and SwapFace, which is the subject of the eighth-most active thread on the forum. In terms of output fidelity, some members praise the tools’ potential to produce high quality deepfakes for simple videos, while others lament their inability to do so for complicated videos. This parallels the striation of the deepfake market into demands for two categories: a lower-tier for “cheapfakes” and a higher-tier for complex scenes that require deeper expertise and extensive image data. The former is handled by these limited-data tools, whose output quality suffers for more complex tasks; buyers of the latter must then turn to the marketplace to pay an expert for their services.
@@
Research Tools.In our annotation of 141 linked Github repositories, we identified that 60 repositories (including forked repositories) were directly linked to 43 academic papers; in aggregate, these open source research code repositories were referenced in the forum over 400 times. The associated papers range in publication date: the oldest is a 2001 paper that achieves color correction between source and destination images using statistical analysis [reinhard2001image], and the six most recent papers were published in 2023, the year we collected our forum data [liu2023dfl, radford2023robust, bain2023whisperx, an2023pano, pan2023drag, yuan2023reliableswap]. Of these papers, CVPR is by far the most prevalent venue, accounting for 13 of the 43 papers (30.2%); the remaining 30 papers are spread across 21 other venues, including well-known conferences ICCV, ECCV, NeurIPS, ICML, and SIGGRAPH. The frequent and up-to-date references to reputable academic publications demonstrates that the MrDeepFakes community follows cutting-edge research.
The For the community favorite, DeepFaceLab DFL (the Github repository was linked 203 times), the corresponding paper was initially published as an ArXiv preprint in 2020 [perov2021dfl] and later published in the peer-reviewed journal Pattern Recognition in 2023. The delta between the two versions largely focuses on potential abuses of DeepFaceLabDFL, which both versions claim is most commonly used by VFX artists. In addressing the ethical issues of advancing deepfake technology, the authors “admit that deepfake techniques may affect public discourse quality and infringe upon the citizens’ portrait rights” when weaponized for “misinformation, manipulation, harassment, [or] persuasion.” Despite this, the authors argue that suppressing the publication would not stop its development and believe that it was their “responsibility to publish DeepFaceLab to the academia community formally.”” [liu2023dfl]. Neither version directly addresses the use or abuse of deepfake technology to generate intimate sexual imagery. In August 2024, the repository for DeepFaceLab DFL was disabled by Github for ToS GitHub for terms of service (ToS) violations.